Perform the indicated operation; write answer in simplest form:
8 and 3/4 – 5 and 1/3


Answer by stanbon(75887)

write answer in simplest form:
8 and 3/4 – 5 and 1/3
----
= (35/4) - (16/3)
lcd = 12
= (105/12) - (64/12)
= 41/12
= 3 5/12
============
Cheers,
Stan H.